# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v7

@@ 4.1.1 Datasets
In our experiments, we extensively utilize four well-established datasets in constrained scenarios for AU detection, namely BP4D [23], DISFA [24], BP4D+ [25], GFT [26], and one well-established dataset in unconstrained scenarios, namely Aff-Wild2 [27]. Our primary focus is on assessing frame-level AU detection, as opposed to datasets that solely provide video-level annotations.
BP4D [23]: This dataset includes 41 participants (23 females, 18 males), each appearing in 8 sessions with 2D and 3D video recordings. It contains approximately 140,000 frames, each labeled for AU occurrence and annotated with 68 facial landmarks. A three-fold cross-validation is used for evaluation.
DISFA [24]: Comprising 27 videos from 15 men and 12 women, this dataset includes 130,788 frames annotated with 68 facial landmarks and AU intensities (0-5). Frames with intensities $\geq$ 2 are positive; others are negative. Three-fold cross-validation is applied for evaluation, similar to BP4D.
BP4D+ [25]: Featuring 82 females and 58 males across 10 sessions, this dataset expands upon BP4D in scale and diversity. Of the 10 sessions, 4 provide 197,875 AU-labeled frames annotated with 49 facial landmarks. Following [4], we train on BP4D and test on BP4D+.
GFT [26]: This dataset records 96 participants in 32 unscripted conversational groups, primarily capturing moderate out-of-plane poses. Annotations include 10 AUs and 68 facial landmarks. Training involves 78 participants (108,000 frames), while testing uses 18 participants (24,600 frames) based on standard splits.
Aff-Wild2 [27]: This in-the-wild dataset includes 305 training videos (1.39 million frames) and 105 validation videos (440,000 frames), sourced from YouTube with diverse demographics and conditions. Each frame is annotated with 12 AUs and 68 facial landmarks. We follow [28] by training on the training set and evaluating on the validation set.
@@
BP4D [23] 21.1 17.1 20.3 46.2 54.9 - 59.4 56.2 46.6 16.9 34.4 16.5 15.2 - -
BP4D+ [25] 10.0 8.0 6.0 50.0 66.0 - 65.0 58.0 60.0 11.0 13.0 17.0 4.0 - -
DISFA [24] 5.0 4.0 15.0 8.1 - 4.3 - 13.2 - - - - - 27.8 8.9
GFT [26] 3.7 13.5 3.7 28.3 - - 24.6 29.3 3.1 10.7 - 25.0 14.1 - -
Aff-Wild2 [27] 11.9 5.1 16.0 26.5 39.9 - 34.5 24.3 - 2.8 - 3.1 2.8 62.8 7.6
@@ 4.1.2 Implementation Details
Following the preprocessing method used in MEGraph [22], the 68 facial landmarks undergo a conversion into 49 facial internal landmarks, excluding those associated with the facial contour. Subsequently, the images undergo augmentation through similarity transformation, encompassing in-plane rotation, uniform scaling, and translation, utilizing the 49 facial landmarks. The resulting images are standardized to a resolution of $\times256$$\times$3 pixels and then randomly cropped into $\times224$$\times$3 pixels, with the addition of a random horizontal flip.
For the pre-training process, the encoder is conducted on one NVIDIA GeForce RTX 3090 GPU with 24GB for one day and implemented by PyTorch using an AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999 and weight decay of $10^{-6}$. We set a learning rate of 10-5 and 60 epochs. Our random seed value is set to 3407, consistent with the linear evaluation protocol. In our discriminative contrastive learning framework, where sample pairs are selected within the same batch, the batch size significantly impacts the experimental results. For consistency, we have set the batch size to 32 in our experiments. In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 12 for BP4D, BP4D+, Aff-Wild2, 10 for GFT, and 8 for DISFA respectively, and the temperature coefficient $\tau$ is 0.5. The analysis for $\beta$ and $p_{1}$, $p_{2}$, $p_{3}$ will introduce in Section. 4.4.2. Additionally, data partitioning is performed using 3-fold cross-validation, where the partitioning is subject-independent to avoid identity leakage.
For the linear evaluation protocol [7, 11], we extract $n_{au}$ 2048-dimensional feature vectors from a model pre-trained with the AUNCE loss. A linear classifier is trained using the AdamW optimizer with the same settings as the pre-training phase, including a batch size of 256, a learning rate of 10-3, and 100 epochs. Results are obtained with weighted cross-entropy loss (WCE).
Although formal significance tests (paired t-test, Wilcoxon) are important, most baselines report only average performance without multiple independent runs. To address this, we report standard deviations (computed over at least five runs) in Tables 3–7, showing the robustness of our framework. Our supplementary analysis indicates that in most experimental scenarios, our method’s performance improvement over baselines exceeds twice the standard deviation.
@@ 4.1.3 Evaluation Metrics
We evaluate our AU detection method using three metrics: F1-macro, F1-micro, and Accuracy. The F1-macro score averages the F1 scores for each AU class. The F1-micro score, as proposed by Hinduja et al. [29], aggregates true positives, false positives, and false negatives across all classes, making it suitable for imbalanced datasets. Accuracy is defined as the ratio of correctly classified frames to total frames. We use F1-score to refer to the F1-macro score.
Table 2: Quantitative comparison results on BP4D dataset, where the evaluation metric is F1-score(%). The values following the $\pm$’ symbol denote the standard deviation (SD) of the data.
@@
MMA-Net [32] 2023 PRL 52.5 50.9 58.3 76.3 75.7 83.8 87.9 63.8 48.7 61.7 46.5 54.4 63.4
SEV-Net [33] 2021 CVPR 58.2 50.4 58.3 81.9 73.9 87.7 87.5 61.6 52.6 62.2 44.6 47.6 63.9
ETDF [34] 2025 TAC 54.7 50.8 57.1 78.8 79.6 84.6 88.0 67.0 54.9 62.9 48.6 54.5 65.1
@@
ARL [35] 2022 TAC 45.8 39.8 55.1 75.7 77.2 82.3 86.6 58.8 47.6 62.1 47.4 55.4 61.1
AC2[3] 2025 IJCV 54.2 54.7 56.5 77.0 76.2 84.0 89.0 63.6 54.8 63.6 46.5 54.8 64.6
ESCDA [5] 2026 PR 56.5 44.5 59.6 79.4 77.4 84.7 88.7 64.9 57.0 65.0 52.5 55.6 65.5
@@
AAR [28] 2023 TIP 53.2 47.7 56.7 75.9 79.1 82.9 88.6 60.5 51.5 61.9 51.0 56.8 63.8
MEGraph [22] 2022 IJCAI 52.7 44.3 60.9 79.9 80.1 85.3 89.2 69.4 55.4 64.4 49.8 55.1 65.5
SACL [36] 2025 TAC 57.8 48.8 59.4 79.1 78.8 84.0 88.2 65.2 56.1 63.8 50.8 55.2 65.6
@@
SimCLR [11] 2020 ICML 38.0 36.4 37.2 66.6 64.7 76.2 76.2 51.1 29.8 56.1 27.5 37.7 49.8
MoCo [12] 2020 CVPR 30.8 41.3 42.1 70.2 70.4 78.7 82.5 53.3 25.2 59.1 31.5 34.3 51.6
EmoCo [19] 2021 FG 50.2 44.7 53.9 74.8 76.6 83.7 87.9 61.7 47.6 59.8 46.9 54.6 61.9
CLP [9] 2024 TIP 47.7 50.9 49.5 75.8 78.7 80.2 84.1 67.1 52.0 62.7 45.7 54.8 62.4
KSRL [13] 2022 CVPR 53.3 47.4 56.2 79.4 80.7 85.1 89.0 67.4 55.9 61.9 48.5 49.0 64.5
CLEF [14] 2023 ICCV 55.8 46.8 63.3 79.5 77.6 83.6 87.8 67.3 55.2 63.5 53.0 57.8 65.9
@@
($\pm$0.15)
($\pm$0.12)
($\pm$0.17)
($\pm$0.21)
($\pm$0.19)
($\pm$0.22)
($\pm$0.22)
($\pm$0.17)
($\pm$0.14)
($\pm$0.15)
($\pm$0.14)
($\pm$0.15)
($\pm$0.17)
Table 3: Quantitative comparison results on DISFA dataset, where the evaluation metric is F1-score(%). The values following the $\pm$’ symbol denote the standard deviation (SD) of the data.
@@
SEV-Net [33] 2021 CVPR 55.3 53.1 61.5 53.6 38.2 71.6 95.7 41.5 58.8
ETDF [34] 2025 TAC 62.6 54.7 70.8 46.3 51.7 76.3 94.4 59.8 64.6
GLEE-Net [37] 2024 TCSVT 61.9 54.0 75.8 45.9 55.7 77.6 92.9 60.0 65.5
@@
ROI [38] 2017 CVPR 41.5 26.4 66.4 50.7 80.5 89.3 88.9 15.6 48.5
ARL [35] 2022 TAC 43.9 42.1 63.6 41.8 40.0 76.2 95.2 66.8 58.7
AC2[3] 2025 IJCV 57.8 59.2 70.1 50.1 54.4 75.1 90.3 66.2 65.4
@@
MEGraph [22] 2022 IJCAI 52.5 45.7 76.1 51.8 46.5 76.1 92.9 57.6 62.4
AAR [28] 2023 TIP 62.4 53.6 71.5 39.0 48.8 76.1 91.3 70.6 64.2
SACL [36] 2025 TAC 62.0 65.7 74.5 53.2 43.1 76.9 95.6 53.1 65.5
@@
SIMCLR [11] 2020 ICML 21.2 23.3 47.5 42.4 35.5 66.8 81.5 52.7 46.4
MoCo [12] 2020 CVPR 22.7 18.2 45.9 45.4 34.1 72.9 83.4 54.5 47.1
CLP [9] 2024 TIP 42.4 38.7 63.5 59.7 38.9 73.0 85.0 58.1 57.4
EmoCo [19] 2021 FG 42.7 41.0 66.3 45.1 50.9 75.5 88.9 58.6 58.6
KSRL [13] 2022 CVPR 60.4 59.2 67.5 52.7 51.5 76.1 91.3 57.7 64.5
CLEF [14] 2023 ICCV 64.3 61.8 68.4 49.0 55.2 72.9 89.9 57.0 64.8
@@
($\pm 0.37$) ($\pm 0.22$) ($\pm 0.42$) ($\pm 0.31$) ($\pm 0.36$) ($\pm 0.49$) ($\pm 0.42$) ($\pm 0.28$) ($\pm 0.36$)
Table 4: Quantitative comparison results on BP4D+ dataset in terms of cross-dataset evaluation, where the evaluation metric is F1-score(%). The values following the $\pm$’ symbol denote the standard deviation (SD) of the data.
@@
AC2[3] 2025 IJCV 42.3 35.4 26.7 80.7 87.0 90.9 85.8 73.3 45.3 43.4 50.3 29.0 57.5
GLEE-Net [37] 2024 TCSVT 39.8 37.9 41.6 83.4 88.2 90.2 87.4 76.6 48.3 42.9 47.7 29.8 59.5
@@
($\pm$ 0.25)
($\pm$ 0.37)
($\pm$ 0.28)
($\pm$ 0.33)
($\pm$ 0.12)
($\pm$ 0.47)
($\pm$ 0.31)
($\pm$ 0.17)
($\pm$ 0.08)
($\pm$ 0.42)
($\pm$ 0.39)
($\pm$ 0.32)
($\pm$ 0.29)
Table 5: Quantitative comparison results on GFT dataset, where the evaluation metric is F1-score(%). The values following the $\pm$’ symbol denote the standard deviation (SD) of the data.
@@
MoCo [12] 2020 CVPR 35.9 45.4 13.5 83.4 71.3 78.1 23.3 37.3 26.6 50.7 46.5
SimCLR [11] 2020 ICML 39.6 48.3 5.6 80.7 76.2 80.6 18.1 41.6 46.1 43.8 48.1
ARL [35] 2022 TAC 51.9 45.9 13.7 79.2 75.5 82.8 0.1 44.9 59.2 47.5 50.1
@@
CLP [9] 2024 TIP 44.6 58.7 34.7 75.9 78.6 86.6 20.3 44.8 56.4 42.2 54.3
AAR [28] 2023 TIP 66.3 53.9 23.7 81.5 73.6 84.2 43.8 53.8 58.2 46.5 58.5
MAL [30] 2023 TAC 52.4 57.0 54.1 74.5 78.0 84.9 43.1 47.7 54.4 51.9 59.8
AC2[3] 2025 IJCV 60.9 58.2 24.4 83.3 75.9 87.4 56.4 46.5 58.3 50.9 60.2
@@
($\pm 0.33$) ($\pm 0.18$) ($\pm 0.28$) ($\pm 0.24$) ($\pm 0.41$) ($\pm 0.11$) ($\pm 0.35$) ($\pm 0.23$) ($\pm 0.42$) ($\pm 0.35$) ($\pm 0.29$)
Table 6: Quantitative comparison results on Aff-Wild2 dataset, where the evaluation metric is F1-score(%). The values following the $\pm$’ symbol denote the standard deviation (SD) of the data.
@@
AAR [28] 2023 TIP 65.4 57.9 59.9 73.2 84.6 83.2 79.9 21.8 27.4 19.9 94.5 41.7 59.1
AC2[3] 2025 IJCV 63.8 53.1 66.0 66.6 80.7 80.1 78.0 30.3 26.5 29.2 93.3 41.4 59.1
@@
($\pm$ 0.25)
($\pm$ 0.37)
($\pm$ 0.28)
($\pm$ 0.33)
($\pm$ 0.12)
($\pm$ 0.47)
($\pm$ 0.31)
($\pm$ 0.17)
($\pm$ 0.08)
($\pm$ 0.42)
($\pm$ 0.39)
($\pm$ 0.32)
($\pm$ 0.29)
[…]
@@
[…]
AC2[3] 2025 IJCV 51.9 49.3 25.8 24.6 50.1 40.9 15.3 20.1 47.2 58.3 38.4
D2CA [20] 2025 TIP 31.0 33.0 43.2 60.5 53.7 71.1 21.1 36.5 36.4 44.6 43.1

### v9

@@ 4.1.1 Datasets
In our experiments, we extensively utilize four well-established datasets in constrained scenarios for AU detection, namely BP4D [zhang:et.al:2014:IVC], DISFA [mavadati:et.al:2013:TAC], BP4D+ [zhang2016multimodal], GFT [girard2017sayette], and one well-established dataset in unconstrained scenarios, namely Aff-Wild2 [kollias2019expression]. Our primary focus is on assessing frame-level AU detection, as opposed to datasets that solely provide video-level annotations.
BP4D [zhang:et.al:2014:IVC]: This dataset includes 41 participants (23 females, 18 males), each appearing in 8 sessions with 2D and 3D video recordings. It contains approximately 140,000 frames, each labeled for AU occurrence and annotated with 68 facial landmarks. A three-fold cross-validation is used for evaluation.
DISFA [mavadati:et.al:2013:TAC]: Comprising 27 videos from 15 men and 12 women, this dataset includes 130,788 frames annotated with 68 facial landmarks and AU intensities (0-5). Frames with intensities $\geq$ 2 are positive; others are negative. Three-fold cross-validation is applied for evaluation, similar to BP4D.
BP4D+ [zhang2016multimodal]: Featuring 82 females and 58 males across 10 sessions, this dataset expands upon BP4D in scale and diversity. Of the 10 sessions, 4 provide 197,875 AU-labeled frames annotated with 49 facial landmarks. Following [shao2021jaa], we train on BP4D and test on BP4D+.
GFT [girard2017sayette]: This dataset records 96 participants in 32 unscripted conversational groups, primarily capturing moderate out-of-plane poses. Annotations include 10 AUs and 68 facial landmarks. Training involves 78 participants (108,000 frames), while testing uses 18 participants (24,600 frames) based on standard splits.
Aff-Wild2 [kollias2019expression]: This in-the-wild dataset includes 305 training videos (1.39 million frames) and 105 validation videos (440,000 frames), sourced from YouTube with diverse demographics and conditions. Each frame is annotated with 12 AUs and 68 facial landmarks. We follow [shao2023facial] by training on the training set and evaluating on the validation set.
@@
BP4D [zhang:et.al:2014:IVC] 21.1 17.1 20.3 46.2 54.9 - 59.4 56.2 46.6 16.9 34.4 16.5 15.2 - -
BP4D+ [zhang2016multimodal] 10.0 8.0 6.0 50.0 66.0 - 65.0 58.0 60.0 11.0 13.0 17.0 4.0 - -
DISFA [mavadati:et.al:2013:TAC] 5.0 4.0 15.0 8.1 - 4.3 - 13.2 - - - - - 27.8 8.9
GFT [girard2017sayette] 3.7 13.5 3.7 28.3 - - 24.6 29.3 3.1 10.7 - 25.0 14.1 - -
Aff-Wild2 [kollias2019expression] 11.9 5.1 16.0 26.5 39.9 - 34.5 24.3 - 2.8 - 3.1 2.8 62.8 7.6
@@ 4.1.2 Implementation Details
Following the preprocessing method used in MEGraph [luo2022learning], the 68 facial landmarks undergo a conversion into 49 facial internal landmarks, excluding those associated with the facial contour. Subsequently, the images undergo augmentation through similarity transformation, encompassing in-plane rotation, uniform scaling, and translation, utilizing the 49 facial landmarks. The resulting images are standardized to a resolution of $\times256$$\times$3 pixels and then randomly cropped into $\times224$$\times$3 pixels, with the addition of a random horizontal flip.
For the pre-training process, the encoder is conducted on one NVIDIA GeForce RTX 3090 GPU with 24GB for one day and implemented by PyTorch using an AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999 and weight decay of $10^{-6}$. We set a learning rate of 10-5 and 60 epochs. In our discriminative contrastive learning framework, where sample pairs are selected within the same batch, the batch size significantly impacts the experimental results. For consistency, we have set the batch size to 32 in our experiments. In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 12 for BP4D, BP4D+, Aff-Wild2, 10 for GFT, and 8 for DISFA respectively, and the temperature coefficient $\tau$ is 0.5. The analysis for $\beta$ and $p_{1}$, $p_{2}$, $p_{3}$ will introduce in Section. 4.4.1.
For the linear evaluation protocol [henaff2020data, chen2020simple], we extract $n_{au}$ 2048-dimensional feature vectors from the model pre-trained using the AUNCE loss. A linear classifier is then trained on these features using AdamW optimizer with the same settings as the pre-training phase, a batch size of 256 on a single GPU, a learning rate of 10-3, and 100 epochs. The results are obtained through training with the weighted cross-entropy loss (WCE).
In addition, we report the standard deviations of the experimental results in Tables 3 through 7. These deviations are computed over at least five independent runs to demonstrate the robustness of the proposed framework.
@@ 4.1.3 Evaluation Metrics
Aligned with current state-of-the-art methodologies, we evaluate our AU detection method using three frame-based metrics: F1-macro score, F1-micro score, and Accuracy. The F1-macro score is obtained by calculating the F1 score for each AU class individually and then averaging these values, which is a standard metric employed by many existing methods. Inspired by the work of Hinduja et al. [hinduja2024time], we also incorporate the F1-micro score, which is computed by aggregating the true positives, false positives, and false negatives across all classes before calculating the overall F1-score. This metric is particularly useful for evaluating performance in imbalanced datasets. Additionally, we include Accuracy, which is defined as the ratio of correctly classified frames to the total number of frames, to provide a more comprehensive assessment of model performance. For simplicity, we will refer to the F1-macro score as F1-score in the remainder of the text.
Table 2: Quantitative comparison results on BP4D dataset, where the evaluation metric is F1-score(%).
@@
MMA-Net [shang2023mma] 2023 PRL 52.5 50.9 58.3 76.3 75.7 83.8 87.9 63.8 48.7 61.7 46.5 54.4 63.4
SEV-Net [yang2021exploiting] 2021 CVPR 58.2 50.4 58.3 81.9 73.9 87.7 87.5 61.6 52.6 62.2 44.6 47.6 63.9
ETDF [chang2025facial] 2025 TAC 54.7 50.8 57.1 78.8 79.6 84.6 88.0 67.0 54.9 62.9 48.6 54.5 65.1
@@
ARL [shao2022facial] 2022 TAC 45.8 39.8 55.1 75.7 77.2 82.3 86.6 58.8 47.6 62.1 47.4 55.4 61.1
AC2[shao2025facial] 2025 IJCV 54.2 54.7 56.5 77.0 76.2 84.0 89.0 63.6 54.8 63.6 46.5 54.8 64.6
ESCDA [shao2026constrained] 2026 PR 56.5 44.5 59.6 79.4 77.4 84.7 88.7 64.9 57.0 65.0 52.5 55.6 65.5
@@
AAR [shao2023facial] 2023 TIP 53.2 47.7 56.7 75.9 79.1 82.9 88.6 60.5 51.5 61.9 51.0 56.8 63.8
MEGraph [luo2022learning] 2022 IJCAI 52.7 44.3 60.9 79.9 80.1 85.3 89.2 69.4 55.4 64.4 49.8 55.1 65.5
SACL [liu2025multi] 2025 TAC 57.8 48.8 59.4 79.1 78.8 84.0 88.2 65.2 56.1 63.8 50.8 55.2 65.6
@@
SimCLR [chen2020simple] 2020 ICML 38.0 36.4 37.2 66.6 64.7 76.2 76.2 51.1 29.8 56.1 27.5 37.7 49.8
MoCo [he2020momentum] 2020 CVPR 30.8 41.3 42.1 70.2 70.4 78.7 82.5 53.3 25.2 59.1 31.5 34.3 51.6
EmoCo [sun2021emotion] 2021 FG 50.2 44.7 53.9 74.8 76.6 83.7 87.9 61.7 47.6 59.8 46.9 54.6 61.9
CLP [li2024contrastive] 2024 TIP 47.7 50.9 49.5 75.8 78.7 80.2 84.1 67.1 52.0 62.7 45.7 54.8 62.4
KSRL [chang2022knowledge] 2022 CVPR 53.3 47.4 56.2 79.4 80.7 85.1 89.0 67.4 55.9 61.9 48.5 49.0 64.5
CLEF [zhang2023weakly] 2023 ICCV 55.8 46.8 63.3 79.5 77.6 83.6 87.8 67.3 55.2 63.5 53.0 57.8 65.9
@@
$\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ ($\pm$
0.15) 0.12) 0.17) 0.21) 0.19) 0.22) 0.22) 0.17) 0.14) 0.15) 0.14) 0.15) 0.17)
Table 3: Quantitative comparison results on DISFA dataset, where the evaluation metric is F1-score(%).
@@
SEV-Net [yang2021exploiting] 2021 CVPR 55.3 53.1 61.5 53.6 38.2 71.6 95.7 41.5 58.8
ETDF [chang2025facial] 2025 TAC 62.6 54.7 70.8 46.3 51.7 76.3 94.4 59.8 64.6
GLEE-Net [zhang2024detecting] 2024 TCSVT 61.9 54.0 75.8 45.9 55.7 77.6 92.9 60.0 65.5
@@
ROI [li2017action] 2017 CVPR 41.5 26.4 66.4 50.7 80.5 89.3 88.9 15.6 48.5
ARL [shao2022facial] 2022 TAC 43.9 42.1 63.6 41.8 40.0 76.2 95.2 66.8 58.7
AC2[shao2025facial] 2025 IJCV 57.8 59.2 70.1 50.1 54.4 75.1 90.3 66.2 65.4
@@
MEGraph [luo2022learning] 2022 IJCAI 52.5 45.7 76.1 51.8 46.5 76.1 92.9 57.6 62.4
AAR [shao2023facial] 2023 TIP 62.4 53.6 71.5 39.0 48.8 76.1 91.3 70.6 64.2
SACL [liu2025multi] 2025 TAC 62.0 65.7 74.5 53.2 43.1 76.9 95.6 53.1 65.5
@@
SIMCLR [chen2020simple] 2020 ICML 21.2 23.3 47.5 42.4 35.5 66.8 81.5 52.7 46.4
MoCo [he2020momentum] 2020 CVPR 22.7 18.2 45.9 45.4 34.1 72.9 83.4 54.5 47.1
CLP [li2024contrastive] 2024 TIP 42.4 38.7 63.5 59.7 38.9 73.0 85.0 58.1 57.4
EmoCo [sun2021emotion] 2021 FG 42.7 41.0 66.3 45.1 50.9 75.5 88.9 58.6 58.6
KSRL [chang2022knowledge] 2022 CVPR 60.4 59.2 67.5 52.7 51.5 76.1 91.3 57.7 64.5
CLEF [zhang2023weakly] 2023 ICCV 64.3 61.8 68.4 49.0 55.2 72.9 89.9 57.0 64.8
@@
$\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ ($\pm$
0.37) 0.22) 0.42) 0.31) 0.36) 0.49) 0.42) 0.28) 0.36)
Table 4: Quantitative comparison results on BP4D+ dataset in terms of cross-dataset evaluation, where the evaluation metric is F1-score(%).
@@
AC2[shao2025facial] 2025 IJCV 42.3 35.4 26.7 80.7 87.0 90.9 85.8 73.3 45.3 43.4 50.3 29.0 57.5
GLEE-Net [zhang2024detecting] 2024 TCSVT 39.8 37.9 41.6 83.4 88.2 90.2 87.4 76.6 48.3 42.9 47.7 29.8 59.5
@@
$\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ ($\pm$
0.18) 0.46) 0.14) 0.32) 0.23) 0.17) 0.14) 0.49) 0.33) 0.25) 0.12) 0.24) 0.26)
Table 5: Quantitative comparison results on GFT dataset, where the evaluation metric is F1-score(%).
@@
MoCo [he2020momentum] 2020 CVPR 35.9 45.4 13.5 83.4 71.3 78.1 23.3 37.3 26.6 50.7 46.5
SimCLR [chen2020simple] 2020 ICML 39.6 48.3 5.6 80.7 76.2 80.6 18.1 41.6 46.1 43.8 48.1
ARL [shao2022facial] 2022 TAC 51.9 45.9 13.7 79.2 75.5 82.8 0.1 44.9 59.2 47.5 50.1
@@
CLP [li2024contrastive] 2024 TIP 44.6 58.7 34.7 75.9 78.6 86.6 20.3 44.8 56.4 42.2 54.3
AAR [shao2023facial] 2023 TIP 66.3 53.9 23.7 81.5 73.6 84.2 43.8 53.8 58.2 46.5 58.5
MAL [li2023meta] 2023 TAC 52.4 57.0 54.1 74.5 78.0 84.9 43.1 47.7 54.4 51.9 59.8
AC2[shao2025facial] 2025 IJCV 60.9 58.2 24.4 83.3 75.9 87.4 56.4 46.5 58.3 50.9 60.2
@@
$\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ ($\pm$
0.33) 0.18) 0.28) 0.24) 0.41) 0.11) 0.35) 0.23) 0.42) 0.35) 0.29)
Table 6: Quantitative comparison results on Aff-Wild2 dataset, where the evaluation metric is F1-score(%).
@@
AAR [shao2023facial] 2023 TIP 65.4 57.9 59.9 73.2 84.6 83.2 79.9 21.8 27.4 19.9 94.5 41.7 59.1
AC2[shao2025facial] 2025 IJCV 63.8 53.1 66.0 66.6 80.7 80.1 78.0 30.3 26.5 29.2 93.3 41.4 59.1
@@
$\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ $\pm ($$\pm$ ($\pm$
0.25) 0.37) 0.28) 0.33) 0.12) 0.47) 0.31) 0.17) 0.08) 0.42) 0.39) 0.32) 0.29)
Table 7: Quantitative comparison results for common 10 AUs of cross evaluation between BP4D and GFT. BP4D → GFT denotes training on BP4D and testing on GFT. The evaluation metric is F1-score(%).
@@
GFT → BP4D Ours (WCE) - 49.5 45.5 26.0 20.4 52.3 36.6 17.3 18.6 50.5 54.4 37.1
AC2[shao2025facial] 2025 IJCV 51.9 49.3 25.8 24.6 50.1 40.9 15.3 20.1 47.2 58.3 38.4
D2CA [li2025decoupled] 2025 TIP 31.0 33.0 43.2 60.5 53.7 71.1 21.1 36.5 36.4 44.6 43.1